Manufacturing costs for Davenport Company during 2018 were as follows:

Beginning Finished Goods, 1/1/18 $ 24,900
Beginning Raw Materials, 1/1/18 36,300
Beginning Work in Process, 1/1/18 111,100
Direct Labor for 2018 276,100
Ending Finished Goods, 12/31/18 23,200
Ending Raw Materials, 12/31/18 40,650
Ending Work in Process, 12/31/18 121,400
Material Purchases for 2018 305,000
(including $19,000 of indirect material)

Required:

1. Compute direct material used.

2. Compute applied overhead if the company applies overhead at a rate of 0.83 (83%) of direct labor cost.

3. Compute total manufacturing cost.

4. Compute cost of goods manufactured.

5. Compute cost of goods sold.
Answer #1

1.Direct material used = Beginning Inventory + Purchases – Ending Inventory

= 36,300+(305000-19000)-40,650

= $281,650

2.Applied overhead = Direct labor cost*83%

= 276100*83%

= $229,163

3.Total manufacturing cost = Direct material + direct labor + overheads

= 281650+276100+229163

= $786,913

4.Cost of goods manufactured = Total manufacturing cost + Beginning work in process inventory – ending work in process inventory

= 786913+111,100-121,400

= $776,613

5.COGS = Cost of goods manufactured + Beginning FG Inventory – Ending FG Inventory

= 776,613 + 24900-23200

= $778,313